OpenStudy (anonymous):

Find all z which satisfy (conjugate z)=z^2

OpenStudy (anonymous):

i tried this one... any ideas?

Directrix (directrix):

My work just disappeared. I set (a - bi) = to the square of (a + bi). I had a = a^2 - b^2 and -b = 2ab. Then, I decided the answers might be 0 + 0i, 1 + 0i, and -1 + 0i but that's just a conjecture.

OpenStudy (anonymous):

i didn't check but I got if z = a + bi, a=-1/2, b=sqrt(3)/2 checking now...

OpenStudy (anonymous):

no, doesn't work.

Directrix (directrix):

On the a = a^2, I had that but remembered that all the real parts had to be set equal and the -b^2 is real so a = a^2 - b^2. Then the imaginary parts are set to be equal. so =b = 2ab. Now, I'm thinking that 0 + 1i and 0 - 1i are solutions.

Directrix (directrix):

So far, i'm considering these: 0 + 0i, 1 + 0i, -1 + 0i, and 0 + 1i and 0 - 1i as solutions.

OpenStudy (anonymous):

yea. looks good.. 4 solutions

Directrix (directrix):

I think there's an elegant way to write a general solution but I don't know what it is.